. A company has annt fixed costs of $2,000 per month and variable costs of S2 per unit produced. If the company charges $12 per unit for their product, what is the breakeven annual production quantity? Assume that the profit expected per month is $10,000, how many units should be sold per month? 2. A concrete mixer has been purchased to improve production in IE 2324 class. There are two operators required to operate the concrete mixer at $20/hour, work 8 hours/day and 20 days/month. Utilities required to operate the mixer cost $2,000/month. The production cost of m concrete is S100 per concrete produced. The revenue to be generated from production of 1 m of concrete is $500. How many m of concrete need to be sold per month so that the revenue generated can breakeven with the expenditure? Also, if an expected $10,000 profit is expected monthly, how many m of concrete need to be sold? IE 2324 Restaurant, which sells only pepperoni pizza, the expenses per pizza are: Fixed Costs Variable Costs revenve 0x General Labor $1,390 Flour S0.50 Rent Insurance$200 Water 0.01 Advertising $100 Cheese $3.00 Utilities$450 Pepperoni $2.00 $2,000 Yeast $0.05 How many pizzas does IE 2324 REstaurant need to sell at $10 each to breakeven? If they sold 300 pizzas the first month, did they make a profit or loss? IE2324 class has created a startup and the app is starting to generate revenues! The startup charges a $ 29 fee from a user, to have a license. The fixed cost is $ 1,500 (per month) for technical support, and variable cost of $ 4 per license sold. It is forecasted that the company will sell 500 licenses per month, for the next year How many months will you reach the breakeven of your investment? How many licenses do they need to sell to breakeven? 4. a) b) Estimate the cost of expanding a planned new clinic by 20,000ft. The appropriate capacity exponent is 0.66, and the budget estimate for 200,000ft2 was $15 million 5. Five years ago, when the relevant cost index was 120, a nuclear centrifuge cost $40,000. The centrifuge had a capacity of separating 1500 gallons of ionized solution per hour. Today, it is desired to build a centrifuge with a capacity 4500 gallons per hour, but the cost index now is 300. Assuming a power sizing exponent to reflect economies of scale, x, of 0.75, use the power sizing

Please answer number 2 and 4. I appreciate you!
Answer #1

Problem 2

Monthly operation time required=t=8 hours a day*20 days a month=160

Operation cost per month=OC=20*160=$3200

Utility cost per month=UC=$2000

Fixed Cost, FC=OC+UC=3200+2000=$5200 per month

Price, P=$500 per cubic meter

Variable Cost=VC=Production Cost=$100 per hour

Break Even quantity=BEP=FC/(P-V)=5200/(500-100)=13 cubic meter

Now desired Profit=PR=$10000 per month

Cubic meter of concrete required=(F+PR)/(P-V)=(5200+10000)/(500-100)=38

Problem 4

License Fee (Price), P=$29

Variable cost=VC=$4 per license

Fixed Cost=FC=$1500 per month

a)

Contribution Margin per ticket=CM=P-V=29-4=$25 per ticket

Expected volume,Q=500 licenses

Expected contribution margin per month=CM*Q=25*500=$12500

In this case Investment is Fixed Cost only.

Let n be number of months required to break even. At this time

Time required to break even=FC/n*Contribution Margin per month=1500/12500=0.12 months

b)

Break even Point=BEP=FC/(P-V)=1500/(29-4)= 60

60 licenses are required to breakeven.